# The Landscape of Agentic Reinforcement Learning for LLMs: A Survey

### v1

@@ 3.1 Planning
Planning, the deliberation over a sequence of actions to achieve a goal, constitutes a cornerstone of artificial intelligence, demanding complex reasoning, world knowledge, and adaptability [69]. While initial efforts leveraged the innate capabilities of LLMs through prompting-based methods [70] (e.g., ReAct [71]), these approaches lacked a mechanism for adaptation through experience. RL has emerged as a powerful paradigm to address this gap, enabling agents to refine their planning strategies by learning from environmental feedback. The integration of RL into agent planning manifests in two distinct paradigms, distinguished by whether RL functions as an external guide to a structured planning process or as an internal driver that directly evolves the LLM’s intrinsic planning policy, which we will detail below.
@@
One major paradigm frames RL as an external guide to the planning process, where the LLM’s primary role is to generate potential actions within a structured search framework. Here, RL is not employed to fine-tune the LLM’s generative capabilities directly, but rather to train an auxiliary reward or heuristic function [66]. This learned function then guides a classical search algorithm, such as Monte Carlo Tree Search (MCTS), by evaluating the quality of different planning trajectories. Representative works like RAP [72] and LATS [73] exemplify this method. They leverage an RL-aided model to assess LLM-generated steps, thereby steering the search toward more promising solutions. In this configuration, the LLM acts as a knowledge-rich action proposer, while RL provides the adaptive, evaluative feedback necessary for efficient exploration.
@@
A second, more integrated paradigm positions RL as an internal driver of the agent’s core planning capabilities. This approach casts the LLM directly as a policy model and optimizes its planning behavior through direct environmental interaction. Instead of guiding an external search algorithm, RL-based feedback from trial and error is used to directly refine the LLM’s internal policy for generating plans. This is achieved through methods derived from RLHF, such as leveraging DPO on successful versus failed trajectories as seen in ETO [74], or through lifelong learning frameworks. For instance, VOYAGER [75] iteratively builds and refines a skill library from environmental interaction. This paradigm transforms the LLM from a static generator into an adaptive policy that continuously evolves, enhancing its robustness and autonomy in dynamic environments. AdaPlan and its PilotRL framework [76] leverage global plan-based guidance with progressive RL to enhance LLM agents’ long-horizon planning and execution coordination in text game environments like AFLWorld and TextCraft.
@@ 3.2 Tool Using
RL has emerged as a pivotal methodology for evolving tool-enabled language agents from post-hoc, ReAct-style pipelines to deeply interleaved, multi-turn Tool-Integrated Reasoning (TIR) systems. While early paradigms successfully demonstrated the feasibility of tool invocation, their reliance on SFT or prompt engineering limited agents to mimicking static patterns, lacking the strategic flexibility to adapt to novel scenarios or recover from errors. Agentic RL addresses this by shifting the learning paradigm from imitation to outcome-driven optimization, enabling agents to autonomously discover when, how, and which tools to deploy. This evolution charts a clear trajectory, which we explore in three stages. We begin with (1) early ReAct-style tool calling, then examine (2) modern tool-integrated reasoning (TIR) that deeply embeds tool use within cognitive loops, and finally, discuss the prospective challenge of (3) multi-turn TIR, focusing on temporal credit assignment for robust, long-horizon performance.
@@
Early paradigms for tool invocation predominantly relied on either prompt engineering or SFT to elicit tool-use behaviors. The (I) prompt engineering approach, exemplified by ReAct [71], leveraged few-shot exemplars to guide an LLM to interleave reasoning traces and actions within a "Thought-Action-Observation" cycle, capitalizing on the model’s in-context learning abilities. Going beyond, (II) SFT-based methods were introduced to internalize models’ tool-use capabilities. Frameworks like Toolformer [77] employed a self-supervised objective to teach models where to insert API calls, while others like FireAct [78], AgentTuning [79], Agent-FLAN [80] fine-tuned models on expert-generated or curated datasets of tool-interaction trajectories (e.g., AgentBank [81], APIBank [82]). Although SFT improved the reliability of tool invocation, both of these early approaches are fundamentally constrained by their imitative nature. They train agents to replicate static, pre-defined patterns of tool use, thereby lacking the strategic flexibility to adapt to novel scenarios or recover from unforeseen errors, a limitation that RL-centric approaches directly address by shifting the learning objective from imitation to outcome-driven optimization.
@@
Building on the limitations of purely imitative paradigms, RL-based approaches for tool use shift the objective from replicating fixed patterns to optimizing end-task performance. This transition enables agents to strategically decide when, how, and in what combination to invoke tools, adapting dynamically to novel contexts and unforeseen failures. At the foundation, frameworks such as ToolRL [83] demonstrate that, even when initialized from base models without any imitation traces, RL training can elicit emergent capabilities, e.g., self-correction of faulty code, adaptive adjustment of invocation frequency, and the composition of multiple tools for complex sub-tasks. Subsequently, a recent surge in research has produced works such as OTC-PO [84], ReTool [85], AutoTIR [86], VTool-R1 [87], DeepEyes [88], Pixel-Reasoner [89], Agentic Reasoning [90], ARTIST [91], ToRL [92] and numerous other works [93, 94, 95, 96, 97, 98, 99, 100, 101], which employ RL policies that interleave symbolic computation (e.g., code execution, image editing) with natural-language reasoning within a single rollout. This integrated control loop allows the agent to balance precise, tool-mediated operations with flexible verbal inference, tailoring the reasoning process to the evolving task state. Recent work [102] theoretically proves that TIR fundamentally expands LLM capabilities beyond the “invisible leash” of pure-text RL by introducing deterministic tool-driven state transitions, establishes token-efficiency arguments for feasibility under finite budgets, and proposes Advantage Shaping Policy Optimization (ASPO) to stably guide agentic tool use.
Today, such tool-integrated reasoning is no longer a niche capability but a baseline feature of advanced agentic models. Mature commercial and open-source systems, such as OpenAI’s DeepResearch and o3 [103], Kimi K2 [104], Qwen QwQ-32B [105], Zhipu GLM Z1 [106], Microsoft rStar2-Agent [107] and Meituan LongCat [108], routinely incorporate these RL-honed strategies, underscoring the centrality of outcome-driven optimization in tool-augmented intelligence.
@@
While tool-integrated RL has proven effective for optimizing actions within a single reasoning loop, the primary frontier lies in extending this capability to robust, long-horizon tasks that require multi-turn reasoning. This leap is fundamentally bottlenecked by the challenge of temporal credit assignment [109]. Current RL approaches often depend on sparse, trajectory-level/outcome-based rewards, making it difficult to pinpoint which specific tool invocation in a long, interdependent sequence contributed to success or failure. While nascent research has begun to explore more granular reward schemes, such as turn-level advantage estimation in GiGPO [110] and SpaRL [111], these are still early steps. Consequently, developing more granular credit assignment mechanisms that can accurately guide the agent through complex decision chains without inadvertently punishing useful exploration or promoting reward hacking remains a critical and largely unsolved problem for advancing agentic systems.
@@ 3.3 Memory
Agentic RL transforms memory modules from passive data stores into dynamic, RL-controlled subsystems, deciding what to store, when to retrieve, and how to forget similar to human [112]. This section traces this evolution through four representative phases.
@@
Early systems (e.g., retrieval-augmented generation) treated memory as an external datastore; when RL was employed at all, it solely regulated when to perform queries. Several classic memory systems without RL involvement, such as MemoryBank [113], MemGPT [114], and HippoRAG [115], adopt predefined memory management strategies that specify how to store, integrate, and retrieve information (e.g., storage via vector databases or knowledge graphs; retrieval based on semantic similarity or topological connectivity). Subsequently, RL was incorporated into the memory management pipeline as a functional component. A notable example is the framework proposed in [116], where the RL policy adjusts retrieval behavior through prospective reflection (multi-level summarization) and retrospective reflection (reinforcing retrieval outcomes). Nevertheless, the memory medium itself remained static (e.g., simple vector store or summary buffer), and the agent exerted no control over the write processes. Recently, Memory-R1 [117] introduces a RL-based memory-augmented Agent framework where a Memory Manager learns to perform structured operations (ADD/UPDATE/DELETE/NOOP) via PPO or GRPO based on downstream QA performance, while an Answer Agent employs a Memory Distillation policy over RAG-retrieved entries to reason and answer.
Table 3: An overview of three classic categories of agent memory; works marked with directly employ RL. The list here is not exhaustive, and we refer readers intersted in borader agent memory to [112].
@@
MemoryBank [113] External Store Static memory with predefined storage/retrieval rules
MemGPT [114] External Store OS-like agent with static memory components
HippoRAG [115] External Store Neuro-inspired memory with heuristic access
Prospect [116] RL-Guided Retrieval Uses RL for reflection-driven retrieval adjustment
@@
MEM1 [119] Explicit Token Memory pool managed by RL to enhance context handling
Memory Token [120] Explicit Token Structured memory for reasoning disentanglement
MemoryLLM [121] Latent Token Latent tokens repeatedly integrated and updated
M+ [122] Latent Token Scalable memory tokens for long-context tracking
IMM [123] Latent Token Decouples word representations and latent memory
Memory [124] Latent Token Forget-resistant memory tokens for evolving context
@@
Zep [125] Temporal Graph Temporal knowledge graph enabling structured retrieval
A-MEM [126] Atomic Memory Notes Symbolic atomic memory units; structured storage
G-Memory [127] Hierarchical Graph Multi-level memory graph with topological structure
Mem0 [128] Structured Graph Agent memory with full-stack graph-based design
@@
Subsequent advancements introduced models equipped with explicit, trainable memory controllers, enabling agents to regulate their own memory states (often stored in token form) without relying on fixed, external memory systems. Notably, such memory is commonly instantiated in two forms. The first is (I) explicit tokens, corresponding to human-readable natural language. For example, in MemAgent [118], the agent maintains a natural-language memory pool alongside the LLM, with an RL policy determining, at each segment, which tokens to retain or overwrite, effectively compressing long-context inputs into concise, informative summaries. Similar approaches include MEM1 [119] and Memory Token [120], both of which explicitly preserve a pool of natural-language memory representations. The second form is (II) implicit tokens, where memory is maintained in the form of latent embeddings. A representative line of work includes MemoryLLM [121] and M+ [122], in which a fixed set of latent tokens serves as “memory tokens.” As the context evolves, these tokens are repeatedly retrieved, integrated into the LLM’s forward computation, and updated, thereby preserving contextual information and exhibiting strong resistance to forgetting. Unlike explicit tokens, these memory tokens are not tied to human-readable text but rather constitute a machine-native form of memory. Related efforts include IMM [123] and Memory [124]. Across both paradigms, these approaches empower agents to autonomously manage their memory banks, delivering significant improvements in long-context understanding, continual adaptation, and self-improvement.
@@
Building on token-level approaches, recent trends are moving toward structured memory representations, which organize and encode information beyond flat token sequences. Representative examples include the temporal knowledge graph in Zep [125], the atomic memory notes in A-MEM [126], and the hierarchical graph-based memory designs in G-Memory [127] and Mem0 [128]. These systems capture richer relational, temporal, or hierarchical dependencies, enabling more precise retrieval and reasoning. However, their management, spanning insertion, deletion, abstraction, and linkage updates, has thus far been governed by handcrafted rules or heuristic strategies. To date, little work has explored the use of RL to dynamically control the construction, refinement, or evolution of such structured memory, making this an open and promising direction for advancing agentic memory capabilities.

### v4

@@ 3.1 Planning
Planning, the deliberation over a sequence of actions to achieve a goal, constitutes a cornerstone of artificial intelligence, demanding complex reasoning, world knowledge, and adaptability (Newell et al., 1958). Initial efforts leveraged the innate capabilities of LLMs through prompting-based methods (Huang et al., 2024a; Yao et al., 2023b). For example, Modular Agentic Planner (MAP) (Webb et al., 2025) introduces a brain-inspired, modular architecture that decomposes planning into specialized LLM modules for conflict monitoring, state evaluation, and coordination. However, these approaches lacked a mechanism for adaptation through experience (Wei et al., 2025a). RL has emerged as a powerful paradigm to address this gap, enabling agents to refine their planning strategies by learning from environmental feedback. The integration of RL into agent planning manifests in two distinct paradigms, distinguished by whether RL functions as an external guide to a structured planning process or as an internal driver that directly evolves the LLM’s intrinsic planning policy, which we will detail below.
@@
One major paradigm frames RL as an external guide to the planning process, where the LLM’s primary role is to generate potential actions within a structured search framework. Here, RL is not employed to fine-tune the LLM’s generative capabilities directly, but rather to train an auxiliary value or heuristic function (Wei et al., 2025a). This learned function then guides a classical search algorithm, such as Monte Carlo Tree Search (MCTS), by evaluating the quality of different planning trajectories. Representative works like RAP (Hao et al., 2023) and LATS (Zhou et al., 2024a) exemplify this approach. Planning without Search (Hong et al., 2025d) extends this idea by leveraging offline goal-conditioned RL to learn a language-based value critic that guides LLM reasoning and planning without updating the LLM’s parameters. In this configuration, the LLM acts as a knowledge-rich action proposer, while RL provides adaptive, evaluative feedback for efficient exploration. Beyond static guidance, Learning When to Plan (Paglieri et al., 2025) formulates dynamic planning as an RL-driven test-time compute allocation problem, training agents to decide when to invoke explicit planning to balance reasoning performance against computational cost. Conversely, MAPF-DT (Atasever et al., 2025) explores the reverse direction, employing Decision Transformer–based offline RL for decentralized multi-agent path planning, with LLM guidance enhancing adaptability and long-horizon efficiency in dynamic environments.
@@
A second, more integrated paradigm positions RL as an internal driver of the agent’s core planning capabilities. This approach casts the LLM directly as a policy model and optimizes its planning behavior through direct environmental interaction. Instead of guiding an external search algorithm, RL-based feedback from trial and error is used to directly refine the LLM’s internal policy for generating plans. This is achieved through methods derived from RLHF, such as leveraging DPO on successful versus failed trajectories as seen in ETO (Song et al., 2024b), or through lifelong learning frameworks. For instance, VOYAGER (Wang et al., 2024a) iteratively builds and refines a skill library from environmental interaction. This paradigm transforms the LLM from a static generator into an adaptive policy that continuously evolves, enhancing its robustness and autonomy in dynamic environments. In a complementary direction, Dynamic Speculative Planning (DSP) (Guan et al., 2025b) embodies an online reinforcement mechanism that adapts the agent’s policy to jointly optimize latency and operational cost, demonstrating that internal policy refinement can govern not only task success but also system efficiency. RLTR (Li et al., 2025p) decouples planning from answer generation and introduces tool-use rewards that directly evaluate action sequence quality, enabling focused optimization of the agent’s planning capability without relying on verifiable final answers. AdaPlan and its PilotRL framework (Lu et al., 2025c) leverage global plan-based guidance with progressive RL to enhance LLM agents’ long-horizon planning and execution coordination in text game environments like AFLWorld and TextCraft. Planner-R1 (Zhu et al., 2025d) examines reward-density effects in Agentic RL, showing that shaped, process-level rewards markedly improve learning efficiency and enable smaller models to attain competitive planning capability.
@@ 3.2 Tool Using
RL has emerged as a pivotal methodology for evolving tool-enabled language agents from post-hoc, ReAct-style pipelines to deeply interleaved, multi-turn Tool-Integrated Reasoning (TIR) systems. While early paradigms successfully demonstrated the feasibility of tool invocation, their reliance on SFT or prompt engineering limited agents to mimicking static patterns, lacking the strategic flexibility to adapt to novel scenarios or recover from errors (Chen et al., 2024g; Kavathekar et al., 2025). Agentic RL addresses this by shifting the learning paradigm from imitation to outcome-driven optimization, enabling agents to autonomously discover when, how, and which tools to deploy. This evolution charts a clear trajectory, which we explore in three stages. We begin with (1) early ReAct-style tool calling, then examine (2) modern tool-integrated reasoning (TIR) that deeply embeds tool use within cognitive loops, and finally, discuss the prospective challenge of (3) multi-turn TIR, focusing on temporal credit assignment for robust, long-horizon performance.
@@
Early paradigms for tool invocation predominantly relied on either prompt engineering or SFT to elicit tool-use behaviors. The (I) prompt engineering approach, exemplified by ReAct (Yao et al., 2023b), leveraged few-shot exemplars to guide an LLM to interleave reasoning traces and actions within a "Thought-Action-Observation" cycle, capitalizing on the model’s in-context learning abilities. Going beyond, (II) SFT-based methods were introduced to internalize models’ tool-use capabilities. Frameworks like Toolformer (Schick et al., 2023) employed a self-supervised objective to teach models where to insert API calls, while others like FireAct (Chen et al., 2023), AgentTuning (Zeng et al., 2024), Agent-FLAN (Chen et al., 2024h) fine-tuned models on expert-generated or curated datasets of tool-interaction trajectories (e.g., AgentBank (Song et al., 2024a), APIBank (Li et al., 2023b)). Although SFT improved the reliability of tool invocation, both of these early approaches are fundamentally constrained by their imitative nature. They train agents to replicate static, pre-defined patterns of tool use, thereby lacking the strategic flexibility to adapt to novel scenarios or recover from unforeseen errors, a limitation that RL-centric approaches directly address by shifting the learning objective from imitation to outcome-driven optimization.
@@
Building on the limitations of purely imitative paradigms, RL-based approaches for tool use shift the objective from replicating fixed patterns to optimizing end-task performance. This transition enables agents to strategically decide when, how, and in what combination to invoke tools, adapting dynamically to novel contexts and unforeseen failures. At the foundation, frameworks such as ToolRL (Qian et al., 2025) demonstrate that, even when initialized from base models without any imitation traces, RL training can elicit emergent capabilities, e.g., self-correction of faulty code, adaptive adjustment of invocation frequency, and the composition of multiple tools for complex sub-tasks. Subsequently, a recent surge in research has produced works such as OTC-PO (Wang et al., 2025e), ReTool (Feng et al., 2025a), AutoTIR (Wei et al., 2025c), VTool-R1 (Wu et al., 2025g), DeepEyes (Zheng et al., 2025g), Pixel-Reasoner (Su et al., 2025a), Agentic Reasoning (Wu et al., 2025e), ARTIST (Singh et al., 2025), ToRL (Li et al., 2025l) and numerous other works (Hao et al., 2025a; Feng et al., 2024a; Wei et al., 2025f; Li et al., 2025f; Wu et al., 2025a; Li et al., 2025i; Chen et al., 2025d; Song et al., 2025d; Ye et al., 2025a), which employ RL policies that interleave symbolic computation (e.g., code execution, image editing) with natural-language reasoning within a single rollout. This integrated control loop allows the agent to balance precise, tool-mediated operations with flexible verbal inference, tailoring the reasoning process to the evolving task state. Lin & Xu (2025) theoretically proves that TIR fundamentally expands LLM capabilities beyond the “invisible leash” of pure-text RL by introducing deterministic tool-driven state transitions, establishes token-efficiency arguments for feasibility under finite budgets, and proposes Advantage Shaping Policy Optimization (ASPO) to stably guide agentic tool use.
Today, such tool-integrated reasoning is no longer a niche capability but a baseline feature of advanced agentic models. Mature commercial and open-source systems, such as OpenAI’s DeepResearch and o3 (OpenAI, 2025), Kimi K2 (Kimi, 2025), Qwen QwQ-32B (Team, 2025c), Zhipu GLM Z1 (AI, 2025), Microsoft rStar2-Agent (Shang et al., 2025a) and Meituan LongCat (Meituan, 2025), routinely incorporate these RL-honed strategies, underscoring the centrality of outcome-driven optimization in tool-augmented intelligence.
@@
While tool-integrated RL has proven effective for optimizing actions within a single reasoning loop, the primary frontier lies in extending this capability to robust, long-horizon tasks that require multi-turn reasoning (Gao et al., 2025c). This leap is fundamentally bottlenecked by the challenge of temporal credit assignment (Pignatelli et al., 2024). Current RL approaches often depend on sparse, trajectory-level/outcome-based rewards, making it difficult to pinpoint which specific tool invocation in a long, interdependent sequence contributed to success or failure. While nascent research has begun to explore more granular reward schemes, such as turn-level advantage estimation in GiGPO (Feng et al., 2025b) and SpaRL (Wang et al., 2025b), these are still early steps. Consequently, developing more granular credit assignment mechanisms that can accurately guide the agent through complex decision chains without inadvertently punishing useful exploration or promoting reward hacking remains a critical and largely unsolved problem for advancing agentic systems.
@@ 3.3 Memory
Agentic RL transforms memory modules from passive data stores into dynamic, RL‑controlled subsystems, deciding what to store, when to retrieve, and how to forget similar to humans (Wu et al., 2025k). This section traces this evolution through four representative phases.
@@
Early systems (e.g., retrieval-augmented generation) treated memory as an external datastore; when RL was employed at all, it solely regulated when to perform queries. Several classic memory systems without RL involvement, such as MemoryBank (Zhong et al., 2024), MemGPT (Packer et al., 2023), and HippoRAG (Gutiérrez et al., 2024), adopt predefined memory management strategies that specify how to store, integrate, and retrieve information (e.g., storage via vector databases or knowledge graphs; retrieval based on semantic similarity or topological connectivity). Subsequently, RL was incorporated into the memory management pipeline as a functional component. A notable example is the framework proposed in Tan et al. (2025b), where the RL policy adjusts retrieval behavior through prospective reflection (multi-level summarization) and retrospective reflection (reinforcing retrieval outcomes). Nevertheless, the memory medium itself remained static (e.g., simple vector store or summary buffer), and the agent exerted no control over the write processes. Recently, Memory-R1 (Yan et al., 2025b) introduced an RL-based memory-augmented Agent framework where a Memory Manager learns to perform structured operations (ADD/UPDATE/DELETE/NOOP) via PPO or GRPO based on downstream QA performance, while an Answer Agent employs a Memory Distillation policy over RAG-retrieved entries to reason and answer. Follow-up works like Mem-$\alpha$ (Wang et al., 2025t) and Memory-as-action (Zhang et al., 2025v) have also explored RL for training agents into automatic memory managers.
Table 3: An overview of three classic categories of agent memory; works marked with directly employ RL. The list here is not exhaustive, and we refer readers interested in broader agent memory to Wu et al. (2025k). The shaded rows indicate the use of reinforcement learning algorithms.
@@
MemoryBank (Zhong et al., 2024) External Store Static memory with predefined storage/retrieval rules
MemGPT (Packer et al., 2023) External Store OS-like agent with static memory components
HippoRAG (Gutiérrez et al., 2024) External Store Neuro-inspired memory with heuristic access
Prospect (Tan et al., 2025b) RL-guided Retrieval Uses RL for reflection-driven retrieval adjustment
@@
Mem-$\alpha{\dagger}$ (Wang et al., 2025t) RL-guided Retrieval RL-guided agents for memory retrieval
Memory-as-action (Zhang et al., 2025v) RL-guided Management End-to-end training agents for memory management
@@
MEM1 (Zhou et al., 2025g) Explicit Token Memory pool managed by RL to enhance context handling
Memory Token (Jin et al., 2025b) Explicit Token Structured memory for reasoning disentanglement
[…]
MemoryLLM (Wang et al., 2024h) Latent Token Latent tokens repeatedly integrated and updated
M+ (Wang et al., 2025s) Latent Token Scalable memory tokens for long-context tracking
IMM (Orlicki, 2025) Latent Token Decouples word representations and latent memory
Memory (Hongkang Yang et al., 2024) Latent Token Forget-resistant memory tokens for evolving context
[…]
@@
Zep (Rasmussen et al., 2025) Temporal Graph Temporal knowledge graph enabling structured retrieval
A-MEM (Xu et al., 2025d) Atomic Memory Notes Symbolic atomic memory units; structured storage
G-Memory (Zhang et al., 2025d) Hierarchical Graph Multi-level memory graph with topological structure
Mem0 (Chhikara et al., 2025) Structured Graph Agent memory with full-stack graph-based design
@@
Subsequent advancements introduced models equipped with explicit, trainable memory controllers, enabling agents to regulate their own memory states (often stored in token form) without relying on fixed, external memory systems. Notably, such memory is commonly instantiated in two forms. The first is (I) explicit tokens, corresponding to human-readable natural language. For example, in MemAgent (Yu et al., 2025d), the agent maintains a natural-language memory pool alongside the LLM, with an RL policy determining, at each segment, which tokens to retain or overwrite, effectively compressing long-context inputs into concise, informative summaries. Similar approaches include MEM1 (Zhou et al., 2025g) and Memory Token (Jin et al., 2025b), both of which explicitly preserve a pool of natural-language memory representations. More frequently, works like ReSum (Wu et al., 2025i), context folding (Sun et al., 2025c) have also explored RL for context memory management. The second form is (II) implicit tokens, where memory is maintained in the form of latent embeddings. A representative line of work includes MemoryLLM (Wang et al., 2024h) and M+ (Wang et al., 2025s), in which a fixed set of latent tokens serves as “memory tokens.” As the context evolves, these tokens are repeatedly retrieved, integrated into the LLM’s forward computation, and updated, thereby preserving contextual information and exhibiting strong resistance to forgetting. Unlike explicit tokens, these memory tokens are not tied to human-readable text but rather constitute a machine-native form of memory. Related efforts include IMM (Orlicki, 2025) and Memory (Hongkang Yang et al., 2024). Across both paradigms, these approaches empower agents to autonomously manage their memory banks, delivering significant improvements in long-context understanding, continual adaptation, and self-improvement. MemGen (Zhang et al., 2025e) for the first time proposes the paradigm of leveraging latent memory tokens for carrying and generating experiential knowledge, posing promising directions for RL-based latent memory.
@@
Building on token-level approaches, recent trends are moving toward structured memory representations, which organize and encode information beyond flat token sequences. Representative examples include the temporal knowledge graph in Zep (Rasmussen et al., 2025), the atomic memory notes in A-MEM (Xu et al., 2025d), and the hierarchical graph-based memory designs in G-Memory (Zhang et al., 2025d) and Mem0 (Chhikara et al., 2025). These systems capture richer relational, temporal, or hierarchical dependencies, enabling more precise retrieval and reasoning. However, their management, spanning insertion, deletion, abstraction, and linkage updates, has thus far been governed by handcrafted rules or heuristic strategies. To date, little work has explored the use of RL to dynamically control the construction, refinement, or evolution of such structured memory, making this an open and promising direction for advancing agentic memory capabilities.